# Planet formation regulated by galactic-scale interstellar turbulence

### v2

@@ 4.2 Implications
  • Disc sub-structures: A diverse range of substructures have been observed in protoplanetary discs (Andrews, 2020). We have demonstrated that the disc is undergoing constant perturbation by in-falling material. We would therefore expect some discs to exhibit structure due to this process (Kuffmeier et al., 2020). For example, by accreting gas from the ISM with a misaligned angular momentum vector we expect disc warping (Dullemond et al., 2022), which may also produce observable shadows (Marino et al., 2015; Krieger et al., 2024). Both external structures and disc warping are evident in some discs (Walsh et al., 2017), and it seems probable that they originate from late stage in-fall at least in some instances.
  • Accretion outbursts and short accretion times: FU Orionis events are short accretion bursts that last $\sim 100$ years (Hartmann & Kenyon, 1996), named for the archetypal example (Herbig, 1966, 1977). In-fall events are among the earliest explanations for these events. There remain several other possible mechanisms such as stellar encounters (Bonnell & Bastien, 1992; Borchert et al., 2022) and thermal instability (Lin et al., 1985; Clarke & Syer, 1996; Lodato & Clarke, 2004), although extended nebulosity surrounding several stars undergoing outburst strongly supports the role of in-fall in at least some cases (Zurlo et al., 2024; Hales et al., 2024). Outbursting star-disc systems may be more broadly categorised as a subset of systems with short accretions time-scales (Manara et al., 2023; Almendros-Abad et al., 2024). Among such discs, a lifetime problem appears particularly pronounced for discs around some Herbig stars (Grant et al., 2023). The assumptions of our model mean that we cannot directly explain these short accretion time-scales; we impose this time-scale for our model. However, given a fixed time-scale $\tau_{\mathrm{acc}}$, discs with short $\tau_{\mathrm{acc}}$ may be episodically replenished in our model. Thus, to a degree the inferred accretion time-scale is decoupled from the expected frequency of observed discs, which no longer needs to be $\sim\tau_{\mathrm{acc}}/\tau_{\mathrm{disc}}$, for disc age $\tau_{\mathrm{disc}}$. Instead, the frequency of discs with a given $\tau_{\mathrm{acc}}$ would be dictated by the frequency of in-fall events that lead to rapid mass gain. Indeed, in our model a small number of discs with very short accretion time-scales ($\sim 0.01$ Myr) are retained for $\sim 3$ Myr (Figure 5). Whether the in-fall itself can also produce a short-lived dip in the accretion time-scale (i.e. an outburst) depends on the turbulence imparted onto the disc, and may depend on the rate or physics of interstellar gas accretion (as discussed above).
@@
  • Planet properties: Possibly the most exciting consequence of protoplanetary discs that are regulated by BHL accretion would be the potential consequences for planets. Multiple studies have shown that mature planet properties are dependent on stellar kinematics (Hamer & Schlaufman, 2019; Winter et al., 2020) Some of these differences could be related to covariance with stellar properties, such as age and metallicity (Miyazaki & Masuda, 2023). However, it is not clear that all correlations can be explained by stellar properties alone (Longmore et al., 2021; Winter & Alexander, 2021; Zink et al., 2023). If discs are constantly replenished by the ISM, stars that form at high velocity relative to the local rest frame should accrete less mass from the ISM after the dispersal of their formation cloud. This represents a direct theoretical link between planet formation and stellar kinematics.
@@ 5 Conclusions
    Material in the disc is constantly replenished over its lifetime. We estimate that $20{-}70$ percent of discs are mostly composed of material captured in the most recent half of their lifetime. This estimate is conservative, ignoring processes apart from stellar accretion that might erode the existing disc, such as planet formation or winds.
@@
    The energy influx due to in-falling material onto the disc is sufficient to drive empirically inferred protoplanetary disc turbulence. We expect a broad distribution for $\alpha_{\mathrm{SS}}$, with typical value $\alpha_{\mathrm{SS}}\sim 10^{-3}$, while some discs with much higher $\alpha_{\mathrm{SS}}\sim 0.1$ (as for DM Tau) and lower $\alpha_{\mathrm{SS}}\sim 10^{-5}$ (as for Oph163131) are expected.
If the ISM drives disc evolution, this represents a considerable change of direction for protoplanetary disc theory, which has largely focused on processes operating on an isolated star-disc system. Consequences of this new picture of disc evolution are fundamental and wide-reaching, relating to every stage of planet formation. The inextricable link between planet formation and the ISM underscores a great urgency for future observational and theoretical studies aiming to quantify how gas in-fall during formation influences protoplanetary disc structures, angular momentum transport and the observed exoplanet population.
Finally, a key advantage of our method is that it can be easily adapted to semi-analytic models in order to include the role of BHL accretion on planet formation. This, as well as unbiased observational surveys to test the findings presented in this work, represent two urgent goals for future studies.

### v3

@@ 4.4 Implications
  • Disc sub-structures: A diverse range of substructures have been observed in protoplanetary discs (Andrews, 2020). We have demonstrated that the disc is undergoing constant perturbation by in-falling material. We would therefore expect some discs to exhibit structure due to this process (Kuffmeier et al., 2020). For example, by accreting gas from the ISM with a misaligned angular momentum vector we expect disc warping (Dullemond et al., 2022; Kimmig & Dullemond, 2024), which may also produce observable shadows (Marino et al., 2015; Bohn et al., 2022; Krieger et al., 2024). Both external structures and disc warping are evident in some discs (Walsh et al., 2017), and it seems probable that they originate from late stage in-fall at least in some instances.
  • Accretion outbursts and short accretion times: FU Orionis events are short accretion bursts that last $\sim 100$ years (Hartmann & Kenyon, 1996), named for the archetypal example (Herbig, 1966, 1977). In-fall of material has been suggested as the origin of these outbursts (e.g. Vorobyov & Basu, 2005), alongside other mechanisms such as stellar encounters (Bonnell & Bastien, 1992; Borchert et al., 2022) and thermal instability (Lin et al., 1985; Clarke & Syer, 1996; Lodato & Clarke, 2004). It is unclear if in-fall events could produce short rise times, but may indirectly contribute by instigating instabilities in the disc (e.g. Deng et al., 2020; Speedie et al., 2023). Extended nebulosity surrounding several stars undergoing outburst supports the role of in-fall in at least some cases (Zurlo et al., 2024; Hales et al., 2024). Outbursting star-disc systems may also be more broadly categorised as a subset of systems with short accretions time-scales (Manara et al., 2023; Almendros-Abad et al., 2024). Among such discs, a lifetime problem appears particularly pronounced for discs around some Herbig stars (Grant et al., 2023). The assumptions of our model mean that we cannot directly explain these short accretion time-scales; we impose this time-scale for our model. Nonetheless, given a fixed time-scale $\tau_{\mathrm{acc}}$, discs with short $\tau_{\mathrm{acc}}$ may be episodically replenished in our model. Thus, to a degree the inferred accretion time-scale is decoupled from the expected frequency of observed discs, which no longer needs to be $\sim\tau_{\mathrm{acc}}/\tau_{\mathrm{disc}}$, for disc age $\tau_{\mathrm{disc}}$. Instead, the frequency of discs with a given $\tau_{\mathrm{acc}}$ would be dictated by the frequency of in-fall events that lead to rapid mass gain. Indeed, in our model a small number of discs with very short accretion time-scales ($\sim 0.01$ Myr) are retained for $\sim 3$ Myr (Figure 5).
@@
  • Planet properties: Possibly the most exciting consequence of protoplanetary discs that are regulated by BHL accretion would be the potential consequences for planets. Multiple studies have shown that mature planet properties are dependent on stellar kinematics (Hamer & Schlaufman, 2019; Winter et al., 2020). Some of these differences could be related to covariance with stellar properties, such as age and metallicity (Miyazaki & Masuda, 2023). However, it is not clear that all correlations can be explained by stellar properties alone (Longmore et al., 2021; Winter & Alexander, 2021; Zink et al., 2023). If discs are constantly replenished by the ISM, stars that form at high velocity relative to the local rest frame should accrete less mass from the ISM after the dispersal of their formation cloud. This represents a direct theoretical link between planet formation and stellar kinematics.
@@ 5 Conclusions
    Material in the disc is constantly replenished over its lifetime. We estimate that $20{-}70$ percent of discs are mostly composed of material captured in the most recent half of their lifetime. This estimate is conservative in the sense that we include only stellar accretion, and ignore processes such as planet formation or winds that may further erode the existing disc.
@@
    The energy influx due to in-falling material onto the disc is sufficient to drive empirically inferred protoplanetary disc turbulence. If this in-fall produces turbulence in the outer disc, then we expect a broad distribution for $\alpha_{\mathrm{SS}}$, with typical value $\alpha_{\mathrm{SS}}\sim 10^{-3}$, while some discs with much higher $\alpha_{\mathrm{SS}}\sim 0.1$ (as for DM Tau) and lower $\alpha_{\mathrm{SS}}\sim 10^{-5}$ are expected. We highlight that stellar accretion is not necessarily driven by turbulence in the outer disc, and whether in-fall can contribute to global angular momentum transport to the inner disc remains an open question.
Our approach is highly simplified, representing a novel method that can be built upon and adapted to semi-analytic models in order to include the role of BHL accretion in planet formation. Several processes may alter the rate of replenishment in the disc, but in particular stellar feedback may curtail late stage accretion for relatively massive stars ($\gtrsim 2\,M_{\odot}$) and stars born in regions of many OB stars. These feedback processes must be considered in future works.
If the ISM plays an important role in disc evolution, this represents a considerable change of direction for protoplanetary disc theory, which has largely focused on processes operating on an isolated star-disc system. Consequences of this new picture of disc evolution are fundamental and wide-reaching, relating to every stage of planet formation. The inextricable link between planet formation and the ISM underscores the importance of future observational and theoretical studies aiming to quantify how gas in-fall during formation influences protoplanetary disc structures, angular momentum transport and the observed exoplanet population. In particular, deep observational surveys aiming to correlate extended (in-falling) structures with disc properties and/or stellar accretion rates would test our findings, and are an urgent goal for the future.